May 27, 2011

Data saved in /data/luci/20110528

Observers: Marie Rex, Emily Mclinden, Marco Pedani

TO: Geno

Scripts observed saved on lucifer in: /home/observer/AZ/May2011/

Data saved in /data/luci/20110528

7:10pm winds are calmer, opening up

7:30pm starting twilight flats

#46 -- 58 K level 12,000 counts -- 9,000 counts #59 -- 80 H level 16,000 counts -- 98,000 counts

7:44pm starting Emily's lamp flats

7:50pm Lucifer lost communication with the telescope

Marco reset the communication

7:52pm restarting Emily's lamp flats

8:00pm Going to pointing star

8:05pm Emily's first target

8:13pm Trouble finding guide star, going to new script

8:24 pm Starting Science script

Lucifer instrument control GUI froze, but script is still running properly

9:00 pm restarting instrument control GUI, looks good.

1" seeing

Going to Emily's second target

10:00pm Going to Emily's 3rd target

10:25pm Running Peter's H-band script, sn09nr.lucifer.scriptH

Seeing 0.85"

Pointing is off, going to pointing star

10:32pm Re-running sn09nr.lucifer.scriptH

#150 -- 158

10:48pm Script crashed, "offsetting error"

Marco is debugging --cannot explain this error

10:53pm restarting script sn09nr.lucifer.scriptH 0.7--0.9" seeing

# 159 --#174

11:04pm Same error, starting script again

#175 --#180

11:12pm Same error, Marco and Geno are trying to debug, restarting script 0.9" seeing

#181--#189

11:20 pm crash and restart

#190 --#216

11:35pm going to Rex field A1914

#217 --220 acquisition scripts

11:45pm starting J-band spec script

#221 -- 225 Seeing 0.6"--1"

#226 --230 Seeing 0.7--1"

#227--235 Seeing 0.7 --1.5"

#236 -- 240 Seeing .7 --1.3"

#241 --242 Seeing .8"

1:47 am doing J-band flats

#243 --258

1:54 am telluric acquisition

#263 --266

2:05 am telluric J-band

#267 --270

2:10 am telluric K-band

#271 --274

2:17 am telluric acquisition

Fatal MOS failure -- "error setting MOS position"

Marco is trying to contact Germany. We don't have Peter's or Walter's numbers. We left a message on Frank's mobile phone. We emailed Walter, Frank Peter

2:55 "the calvary has arrived"

3:22 And we're back

A1914 has set, going to do MACS 2043 imaging script Preset failed "primary mirror collimation reached the limit"

Geno did pointing check

3:32am resending preset to macs 2043

Seeing 1.1" --2.5"

#275--292

#293--328 1.4" seeing

#328 --340 1.7" seeing

#341 --382 2" seeing

#383 --400 2" seeing

4:32 Running Endof Night Script

4:45 Running DoDarks.txt

- XiaohuiFan - 27 May 2011
Topic revision: r9 - 28 May 2011, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback